Solving Problems Involving Factor of Polynomials (Geometric Relation Problem)
1. The longest side of the triangle is 2 more than the shortest side. The middle side is 1 more than the shortest side. Find the lengths of the side of the triangle. Show your solution.
2. The dimensions of a rectangular flower garden in Olongapo Park are 20m by 10m. There is a paved walk uniform width surrounding the garden. The combined area of the garden and walk is 600m². Find the width of the walk. Show your solution.

Found 3 solutions by mananth, josgarithmetic, ikleyn:
Answer by mananth(16946)


The dimensions of a rectangular flower garden in Olongapo Park are 20m by 10m. There is a paved walk uniform width surrounding the garden. The combined area of the garden and walk is 600m². Find the width of the walk. Show your solution.
let width of walkway be x
The dimensions of a rectangular flower garden in Olongapo Park are 20m by 10m.
The length and width including walkway will be
(20+2x)(10+2x)
The combined area of the garden and walk is 600m².
(20+2x)(10+2x)= 600
200+40x+20x+4x^2=600
4x^2+60x-400=0
divide by 4
x^2+15x-100=0
x^2 +20x-5x-100=0
(x+20)(x-5)=0
x=-20 or x=5
Can width be negative ?
So x = 5
the width of the walk is 5 meters

















Answer by josgarithmetic(39618)

LONGEST                x+2
MIDDLE                 x+1
SHORTEST                x

Not enough to find the lengths of the sides of this triangle. About all you can is the perimeter is 3x+3. You could try using Triangle Inequality Theorem to see if the number do or not work.

Answer by ikleyn(52790)

.

If you want to get full,  complete and correct answers/solutions from tutors
at this forum,  NEVER  pack more than one problem per post.

MEMORIZE :   only one problem/question per post.

It is the  POLICY  and the  RULE  of this forum.

Those who violate this rule work against their own interests.


Come again soon and submit  ONE  problem per post.


Do not work against yourself.